# Overcoming Output Dimension Collapse: How Sparsity Enables Zero-shot Brain-to-Image Reconstruction at Small Data Scales

### v1

@@ 4 Empirical study: ODC on real data
In the previous section, we showed that naive multivariate linear regression traps in output dimension collapse at small data scales. This section examines how this collapse affects brain-to-image reconstruction in practice. Specifically, we first analyze the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ under various data scales and then compare $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the actual prediction from brain $\hat{\bm{y}}_{\mathrm{te}}$.
@@
Our experiments mainly follow the method of Shen et al. (2019), which employs a Translator–Generator framework to reconstruct images from brain activity. The original method adopts a sparse linear regression model as the translator, and the generator iteratively optimizes the reconstructed image via a CNN-based process. Here, we replace the sparse translator with naive multivariate linear regression to observe the effect of output dimension collapse. For the generator, we adopt Deep Image Prior (DIP) (Ulyanov et al., 2018) as the image prior to visualize the latent features with minimal prior influence, following the protocol of Shen et al. (2019) for iterative image optimization. We also evaluated the method of Ozcelik and VanRullen (2023), which adopts naive multivariate linear regression as the translator (Figure 2). We selected the Deeprecon dataset (Shen et al., 2019), designed explicitly for brain-to-image reconstruction. The training set comprises 1,200 natural images from ImageNet (Deng et al., 2009), and the test set consists of 50 natural images from different categories (unseen during training) and 40 artificial shapes drawn entirely from outside the training domain. Zero-shot prediction beyond the training stimuli is crucial in brain-to-image reconstruction research, so we choose a dataset whose training and test partitions are strictly separated, enabling a rigorous zero-shot evaluation.
@@
To see how the data scale drives output dimension collapse, we progressively reduce the training samples from 1,200 down to 600, 300, and 150, and we compute the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ in each condition. Figure 3A plots the best prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ for each test sample across these different values of $n$. As $n$ decreases, the error grows, indicating that the dimension of the training subspace is shrinking and cannot fully capture the variability in the true latent features. Figure 3B shows example reconstructions generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$ and the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ at each data scale. While the reconstruction from the uncollapsed true feature $\mathcal{G}(\bm{y}_{\mathrm{te}})$ is nearly perfect, the quality of the reconstruction from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ worsens noticeably as $n$ decreases. These results suggest that in reconstruction methods using naive multivariate linear regression, the data scale critically affects reconstruction quality through output dimension collapse.
@@
Figure 3: The best predictions for different training data sizes. (A) The best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ for each test sample across training data sizes 1200, 600, 300, and 150. The vertical axis represents the error of the best prediction, and the horizontal axis represents the sample index sorted by the error at n=n=150. (B) Representative reconstructions from the best predictions for different training data sizes. From top: ground truth, reconstructed images generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best prediction values $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ using 1200, 600, 300, and 150 samples. The left four columns show natural images, and the right two columns show artificial shape images.
To assess how output dimension collapse limits actual prediction performance, we use the full set of 1,200 training images and compare the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the prediction from brain activity $\hat{\bm{y}}_{\mathrm{te}}$. Figure 4A compares the prediction errors of the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ and the brain prediction $\hat{\bm{y}}_{\mathrm{te}}$. For natural images, the best prediction error accounts for 70% of the total prediction error, and even for artificial images, the proportion reaches 40% (left panel). Examining individual samples reveals a strong correlation between the best and brain prediction errors (right panel). Samples that lie farther from the subspace (i.e., with larger best prediction errors) exhibit correspondingly larger total prediction errors. These observations show that a substantial portion of the prediction error corresponds to the irreducible out-of-subspace error, i.e., a direct consequence of output dimension collapse. Figure 1 presents the corresponding results with subsampled training sets. Figure 4B shows the reconstructed images generated from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ and the brain prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. Both reconstructions are similarly blurred, markedly different from the “true” reconstruction. We observe similar results with another reconstruction method (Ozcelik and VanRullen, 2023) (Figure 2). As a reference point, Shen et al.’s original method makes different design choices (e.g., a sparse translator) and can yield sharper images.
@@
Figure 4: Comparison of the best prediction and the brain prediction. (A) Left: Percentage of the best prediction error to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Blue for natural images, orange for artificial shapes. The error bars denote the standard deviation across samples. Right: Sample-wise comparison of best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Each dot represents one image, and the dotted line indicates where the two errors are equal. (B) Representative reconstructions generated from the best predictions and brain predictions. From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. The bottom row shows the latest method reconstructions (Shen et al., 2019). The left four columns show natural images, and the right two columns show artificial shape images.
@@
These results confirm that naive multivariate linear regression suffers severely from output dimension collapse due to small data scales. Even the best possible subspace prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ deviates substantially from the true latent $\bm{y}_{\mathrm{te}}$, creating an irreducible error that impedes zero-shot reconstruction. In turn, actual predictions $\hat{\bm{y}}_{\mathrm{te}}$ inherit this limitation and produce blurred reconstructions. These results underscore that naive regression is incompatible with the fundamental structure of reconstruction data: the vast image space and the limited training data.
@@ 5 Leveraging sparse brain-to-feature structure
In the previous sections, we demonstrated that naive multivariate linear regression suffers from output dimension collapse at small data scales, restricting predictions to the low-dimensional subspace spanned by the training outputs, thereby limiting generalization to unseen stimuli. This limitation arises from the data structure of brain-to-image reconstruction: the vast image space and the limited training data. In contrast, brain-to-feature mapping is known to be sparse, and several studies have reported improved prediction performance by exploiting this sparsity. In this section, we demonstrate that sparse linear regression, which leverages this structure, can yield low prediction errors even at small data scales. We first show that sparse regression is not subject to output dimension collapse. We then formalize a data generation process and estimate the prediction error characterized by a selection rate $p_{\mathrm{sel}}$ and a hit rate $\pi$.
@@
In brain-to-image reconstruction, employing “brain-like” latent features renders the brain-to-feature mapping sparse. Neurophysiology shows that the visual cortex exhibits local or selective activity in response to visual stimuli (Hubel and Wiesel, 1962; De Valois et al., 1982; Vinje and Gallant, 2000). Correspondingly, several reconstruction studies have adopted latent representations that exhibit similar locality and selectivity, such as local bases or intermediate CNN features (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019). Under these “brain-like” features, only a small subset of voxels should influence each latent dimension, so the input-output mapping becomes sparse (Figure 5A).
@@
Figure 5: Sparse brain-to-feature mapping. (A) If the brain and the latent features exhibit local or selective activity, their relationship should become sparse mapping. (B) Data generation process that reflects the sparse structure of the brain-to-feature mapping. The smaller the non-zero weight ratio $a$, the sparser the input-output mapping is.
Sparse regression models can effectively exploit the sparse brain-to-feature mapping. The model selects a subset of input variables to assign non-zero weights, thereby reducing the effective number of free parameters. Sparse regression encompasses approaches that first perform variable selection and then fit a model (wrapper methods (Kohavi and John, 1997) and filter methods such as CFS (Hall, 2000) and SIS (Fan and Lv, 2008)), as well as approaches that perform model fitting and variable selection jointly (embedded methods such as Lasso (Tibshirani, 1996) and ARD (MacKay, 1992)). Importantly, these models are not subject to output dimension collapse at small data scales ($n<d_{\mathrm{out}}$). Naive regression uses a single shared set of input variables to predict all output dimensions, which implies $\mathrm{rank}(\hat{W})\leq n$. In contrast, a sparse model can select a different subset of input variables for each output dimension, allowing $\mathrm{rank}(\hat{W})>n$. Conceptually, ODC results from a mismatch between the implicit parameter reduction of the naive model and the data structure. When the true mapping is sparse, explicit reduction through variable selection that matches this structure should be preferred. Theoretically, sparse regression can achieve accurate prediction with small data scales by leveraging sparsity (Wainwright, 2009a; Donoho, 2006; Fan and Lv, 2008). Empirically, reconstruction methods that employ brain-like latent features and sparse regression models have achieved accurate zero-shot reconstruction even with limited training data (Miyawaki et al., 2008; Shen et al., 2019).
@@
To engage in a deeper discussion, we formalize a data generation process that reflects the sparse structure of the brain-to-feature mapping (Figure 5B). Let us assume that the brain activity $\bm{x}\in\mathbb{R}^{d}$ and latent features $\bm{y}\in\mathbb{R}^{d}$ have the same dimensionality $d$ for simplicity. Each sample pair $\left(\bm{x}^{(i)},\bm{y}^{(i)}\right)$ is generated as follows:
@@
with $S_{k}\subseteq\{1,2,\ldots,d\}$ a random subset of size $s$ chosen uniformly at random from the $d$ indices. We denote the non-zero weight ratio by $a=s/d$, reflecting the fraction of input dimensions that contribute to each output dimension. The smaller the non-zero weight ratio $a$, the sparser the input-output mapping is.
Under this generative setting (Figure 5B), we analyze the prediction error of a generic sparse-regression framework trained on the $n$ training pairs $\{(\bm{x}^{(i)},\bm{y}^{(i)})\}_{i=1}^{n}$. The model first selects a subset of the inputs and then fits a ridge estimator with an optimally tuned penalty on that support. Let the selection rate $p_{\mathrm{sel}}\in(0,1]$ be the fraction of the $d$ inputs that are selected, and the hit rate $\pi\in[0,1]$ be the fraction of the $s$ truly relevant inputs that are selected. The post-selection problem is an ordinary ridge regression, so an asymptotic theory of Dicker (2016) applies directly: as the sample size $n$ and the dimensionality $d$ both diverge ($n,d\to\infty$) with the data scale $n/d=\gamma<\infty$, the prediction error $R=\frac{1}{d}\mathbb{E}\bigl[\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}\bigr]$ is given by
@@
(see A.3 for the derivation). Throughout the paper, the relation $\approx$ indicates asymptotic equality as $n,d\to\infty$ with $n/d<\infty$. In case $p_{\mathrm{sel}}=1$ and $\pi=1$ (all $d$ inputs selected), this expression reduces to the prediction error of ridge regression without variable selection. This expression allows us to estimate the prediction error $R$ from the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ for a broad class of sparse regression methods, including wrapper and filter methods. Fully embedded methods such as Lasso (Tibshirani, 1996) or ARD (MacKay, 1992) do not satisfy this framework exactly, but refitting ridge on the selected support aligns them with this framework and is known to yield similar prediction performance (Belloni and Chernozhukov, 2013). Figure 6 shows heatmaps of the prediction error $R$ for sparse regression across combinations of the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$. With $p_{\mathrm{sel}}$ fixed, $R$ decreases as $\pi$ increases. Along the diagonal $p_{\mathrm{sel}}=\pi$ (random selection), the error does not decrease, whereas variable selection that attains a high hit rate with a small $p_{\mathrm{sel}}$ does reduce the error.
Figure 6: Prediction error of sparse regression. Each heatmap shows the prediction error $R$ of sparse regression with data scale $n/d=0.2$ (left) and $n/d=0.8$ (right). The horizontal axis represents selection rate $p_{\mathrm{sel}}$, and the vertical axis represents hit rate $\pi$. The diagonal line indicates the case where $p_{\mathrm{sel}}=\pi$, which corresponds to random selection. The upper-left shaded region defined by $\pi\geq p_{\mathrm{sel}}/a$ represents the $(p_{\mathrm{sel}},\pi)$ values that are, in principle, unattainable when $a=0.08$.
Sparse regression reduces prediction error by leveraging the sparse brain-to-feature mapping, a structure that arises when brain-like latent features are employed. By formalizing this sparse structure, we have shown that appropriate variable selection can lead to accurate prediction even at small data scales.

### v2

@@ 4 Empirical study: ODC on real data
We showed that naive multivariate linear regression is prone to output dimension collapse at small data scales. We next examine how this collapse affects brain-to-image reconstruction in practice. Specifically, we first analyze the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ under various data scales and then compare $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the actual prediction from brain $\hat{\bm{y}}_{\mathrm{te}}$.
@@
Our experiments mainly follow the method of Shen et al. (2019), which employs a Translator–Generator framework to reconstruct images from brain activity (see also Appendix B.3). To observe the effect of ODC, we replace their sparse translator with a naive multivariate linear regression model. For the generator, we adopt Deep Image Prior (DIP) (Ulyanov et al., 2018) as the image prior to visualize the latent features with weak prior influence, following the protocol of Shen et al. (2019) for iterative image optimization. We also evaluated the method of Ozcelik and VanRullen (Ozcelik and VanRullen, 2023) as an additional baseline, which adopts naive multivariate linear regression as the translator. We selected the Deeprecon dataset (Shen et al., 2019), designed explicitly for brain-to-image reconstruction. The training set comprises 1,200 natural images from ImageNet (Deng et al., 2009), and the test set consists of 50 natural images from different categories (unseen during training) and 40 artificial shapes drawn entirely from outside the training domain. Zero-shot prediction beyond the training stimuli is crucial in brain-to-image reconstruction research, so we choose a dataset whose training and test partitions are strictly separated, enabling rigorous zero-shot evaluation.
@@
To see how the data scale drives ODC, we progressively reduce the number of training samples from 1,200 down to 600, 300, and 150, and we compute the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ in each condition. Fig. 3A plots the best prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ for each test sample across these different values of $n$. As $n$ decreases, the error grows, indicating that the dimension of the training subspace is shrinking and cannot fully capture the variability in the true latent features. Fig. 3B shows example reconstructions generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$ and the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ at each data scale. While the reconstruction from the uncollapsed true feature $\mathcal{G}(\bm{y}_{\mathrm{te}})$ is nearly perfect, the quality of the reconstruction from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ worsens noticeably as $n$ decreases. These results suggest that in reconstruction methods using naive multivariate linear regression, the data scale critically affects reconstruction quality through ODC.
@@
Figure 3: The best predictions for different training data sizes. (A) The best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ for each test sample across training data sizes 1200, 600, 300, and 150. The vertical axis represents the error of the best prediction, and the horizontal axis represents the sample index sorted by the error at n=n=150. (B) Representative reconstructions from the best predictions for different training data sizes. From top: ground truth, reconstructed images generated from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ using 1200, 600, 300, and 150 samples. The left four columns show natural images, and the right two columns show artificial shape images.
@@
Figure 4: Comparison of the best prediction and the brain prediction. (A) Left: Percentage of the best prediction error relative to the brain prediction error $\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}/\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Blue for natural images, orange for artificial shapes. The error bars denote the standard deviation across samples. Right: Sample-wise comparison of best prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}-\bm{y}_{\mathrm{te}}\|^{2}$ and brain prediction error $\frac{1}{d_{\mathrm{out}}}\|\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\|^{2}$. Each dot represents one image, and the dotted line indicates where the two errors are equal. (B) Representative reconstructions generated from the best predictions and brain predictions. From top: ground truth, reconstructions from the true features $\mathcal{G}(\bm{y}_{\mathrm{te}})$, the best predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$, brain predictions $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. The bottom row shows reconstructions from the original method of Shen et al. (2019). The left four columns show natural images, and the right two columns show artificial shape images.
To assess how ODC limits actual prediction performance, we use the full set of 1,200 training images and compare the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ with the prediction from brain activity $\hat{\bm{y}}_{\mathrm{te}}$. Fig. 4A compares the prediction errors of the best prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ and the brain prediction $\hat{\bm{y}}_{\mathrm{te}}$. For natural images, the best prediction error accounts for 70% of the brain prediction error, and even for artificial images, the proportion reaches 40% (left panel). Examining individual samples reveals a strong correlation between the best and brain prediction errors (right panel). Samples that lie farther from the subspace (i.e., with larger best prediction errors) exhibit correspondingly larger total prediction errors. These observations show that a substantial portion of the prediction error corresponds to the irreducible out-of-subspace error, i.e., a direct consequence of ODC. Fig. A1 presents the corresponding results with subsampled training sets. Fig. 4B shows the reconstructed images generated from the best prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}})$ and the brain prediction $\mathcal{G}(\hat{\bm{y}}_{\mathrm{te}})$. Both reconstructions are similarly blurred, markedly different from the “true” reconstruction. We observe similar results with another reconstruction method (Ozcelik and VanRullen, 2023) (Fig. A2). As a reference point, Shen et al.’s original method (Shen et al., 2019) makes different design choices (e.g., a sparse translator) and can yield sharper images.
@@
These results confirm that naive multivariate linear regression suffers severely from ODC due to small data scales. Even the best possible subspace prediction $\hat{\bm{y}}_{\mathrm{te}}^{\mathrm{opt}}$ deviates substantially from the true latent $\bm{y}_{\mathrm{te}}$, creating an irreducible error that impedes zero-shot reconstruction. In turn, actual predictions $\hat{\bm{y}}_{\mathrm{te}}$ inherit this limitation and produce blurred reconstructions. These results underscore that naive regression is poorly suited to the fundamental structure of reconstruction datasets, in which a vast stimulus space must be learned from only a limited number of paired brain–image samples.
@@ 5 Leveraging sparse brain-to-feature structure
We have shown that naive multivariate linear regression suffers from output dimension collapse at small data scales, restricting predictions to the low-dimensional subspace spanned by the training outputs, thereby limiting generalization to unseen stimuli. This limitation arises from the data structure of brain-to-image reconstruction: the vast image space and the limited training data. In contrast, the brain-to-feature mapping is commonly assumed to be sparse, and several studies have reported successful prediction performance under such sparsity assumptions (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019). We therefore theoretically analyze sparse regression to clarify how exploiting sparsity affects prediction. We first show that sparse regression is not forced into ODC in the same way as naive regression. We then analyze prediction error in a student–teacher setting for a broad class of sparse regression procedures.
@@
In brain-to-image reconstruction, using “brain-like” latent features makes it reasonable to assume a sparse brain-to-feature mapping. Neurophysiology shows that the visual cortex exhibits local or selective activity in response to visual stimuli (Hubel and Wiesel, 1962; De Valois et al., 1982; Vinje and Gallant, 2000). Correspondingly, several reconstruction studies have adopted latent representations that exhibit similar locality and selectivity, such as local bases or intermediate CNN features (Miyawaki et al., 2008; Zhang et al., 2018; Shen et al., 2019). Under these “brain-like” features, only a small subset of voxels should influence each latent dimension, so the input-output mapping becomes sparse (Fig. 5A).
Sparse regression provides an inductive bias that is well aligned with the brain-to-feature mapping. Sparse regression selects a subset of input variables to assign non-zero weights, thereby reducing the effective number of free parameters. Sparse regression encompasses approaches that first perform variable selection and then fit a model (wrapper methods (Kohavi and John, 1997) and filter methods such as CFS (Hall, 2000) and SIS (Fan and Lv, 2008)), as well as approaches that perform model fitting and variable selection jointly (embedded methods such as Lasso (Tibshirani, 1996) and ARD (MacKay, 1992)).
@@
Figure 5: Sparse brain-to-feature mapping. (A) If the brain and the latent features exhibit local or selective activity, their mapping should become sparse. (B) Teacher model in our student–teacher framework, reflecting the sparse structure of the brain-to-feature mapping. The smaller the non-zero weight ratio $a$, the sparser the input-output mapping is.
[…]
@@
To clarify how the translator behaves under sparsity, we introduce a student–teacher model(Fig. 5B). For simplicity, we assume that the brain activity $\bm{x}\in\mathbb{R}^{d}$ and latent features $\bm{y}\in\mathbb{R}^{d}$ have the same dimensionality $d$. Each sample pair $\left(\bm{x}^{(i)},\bm{y}^{(i)}\right)$ is generated by a linear teacher as
@@
with $S_{k}\subseteq\{1,2,\ldots,d\}$ a random subset of size $s$ chosen uniformly at random from the $d$ indices. We denote the non-zero weight ratio by $a=s/d$, which represents the fraction of input dimensions that contribute to each output dimension; the smaller $a$ is, the sparser the input–output mapping.
Under this student–teacher setting (Fig. 5B), we analyze the prediction error of a generic sparse-regression framework trained on the $n$ training pairs $\{(\bm{x}^{(i)},\bm{y}^{(i)})\}_{i=1}^{n}$. The model first selects a subset of the inputs and then fits a ridge estimator with an optimally tuned penalty on that support. Let the selection rate $p_{\mathrm{sel}}\in(0,1]$ be the fraction of the $d$ inputs that are selected, and the hit rate $\pi\in[0,1]$ be the fraction of the $s$ truly relevant inputs that are selected. The post-selection problem is an ordinary ridge regression, so an asymptotic theory of Dicker (2016) applies directly: as the sample size $n$ and the dimensionality $d$ both diverge ($n,d\to\infty$) with the data scale $n/d=\gamma<\infty$, the prediction error $R=\frac{1}{d}\mathbb{E}\bigl[\lVert\hat{\bm{y}}_{\mathrm{te}}-\bm{y}_{\mathrm{te}}\rVert^{2}\bigr]$ is given by
@@
(see Appendix B.4 for the derivation). Throughout the paper, the relation $\approx$ indicates asymptotic equality as $n,d\to\infty$ with $n/d<\infty$. In the special case $p_{\mathrm{sel}}=1$ and $\pi=1$ (all $d$ inputs selected), this expression reduces to the prediction error of ridge regression without variable selection. This expression allows us to estimate the prediction error $R$ from the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ for a broad class of sparse regression methods, including wrapper and filter methods. Fully embedded methods such as Lasso (Tibshirani, 1996) or ARD (MacKay, 1992) do not satisfy this framework exactly, but refitting the ridge on the selected support aligns them with this framework and is known to yield similar prediction performance (Belloni and Chernozhukov, 2013). Fig. 6 shows heatmaps of the prediction error $R$ for sparse regression across combinations of the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$. With $p_{\mathrm{sel}}$ fixed, $R$ decreases as $\pi$ increases. Compared with the no-selection baseline $(p_{\mathrm{sel}},\pi)=(1,1)$, random selection along the diagonal $p_{\mathrm{sel}}=\pi$ does not reduce the error, whereas variable selection that attains a high hit rate with a small $p_{\mathrm{sel}}$ does reduce the error.
Figure 6: Prediction error of sparse regression in the student–teacher model. Each heatmap shows the prediction error $R$ of sparse regression with data scale $n/d=0.2$ (left) and $n/d=0.8$ (right). The horizontal axis represents selection rate $p_{\mathrm{sel}}$, and the vertical axis represents hit rate $\pi$. The diagonal line indicates the case where $p_{\mathrm{sel}}=\pi$, which corresponds to random selection. The upper-left shaded region defined by $\pi\geq p_{\mathrm{sel}}/a$ represents the $(p_{\mathrm{sel}},\pi)$ values that are, in principle, unattainable when $a=0.08$.
Sparse regression can exploit a sparse brain-to-feature mapping and is not structurally constrained by ODC. Our analysis within the student–teacher framework shows that variable selection, which achieves a high hit rate, can lead to accurate prediction even at small data scales. Our error expression in Eq. 12 applies to a broad class of sparse translators beyond the specific correlation-based filter we study next.